# HiDe-LLaVA: Hierarchical Decoupling for Continual Instruction Tuning of Multimodal Large Language Model

### v1

@@ Abstract
Instruction tuning is widely used to improve a pre-trained Multimodal Large Language Model (MLLM) by training it on curated task-specific datasets, enabling better comprehension of human instructions. However, it is infeasible to collect all possible instruction datasets simultaneously in real-world scenarios. Thus, enabling MLLM with continual instruction tuning is essential for maintaining their adaptability. However, existing methods often trade off memory efficiency for performance gains, significantly compromising overall efficiency. In this paper, we propose a task-specific expansion and task-general fusion framework based on the variations in Centered Kernel Alignment (CKA) similarity across different model layers when trained on diverse datasets. Furthermore, we analyze the information leakage present in the existing benchmark and propose a new and more challenging benchmark to rationally evaluate the performance of different methods. Comprehensive experiments showcase a significant performance improvement of our method compared to existing state-of-the-art methods. Our code will be public available.
@@
Haiyang Guo1,2thanks: Equal Contribution., Fanhu Zeng2,311footnotemark: 1, Ziwei Xiang2,3, Fei Zhu4, Da-Han Wang5, Xu-Yao Zhang2,3, Cheng-Lin Liu2,3, 1SAIS, UCAS 2MAIS, Institute of Automation, Chinese Academy of Sciences 3SAI, UCAS 4Centre for Artificial Intelligence and Robotics, HKISI-CAS 5Department of Computer and Information Engineering, Xiamen University of Technology
@@ 1 Introduction
Recent years have witnessed remarkable advancements in Multimodal Large Language Models (MLLMs) Yin et al. (2023) , which extend the capabilities of Large Language Models Touvron et al. (2023) through sophisticated vision-text feature alignment mechanisms Liu et al. (2024b) and autoregressive generation frameworks. The integration of large-scale training corpora and extensive model parameters Yang et al. (2024); Kaplan et al. (2020) has enabled MLLMs to achieve state-of-the-art performance across diverse downstream applications Zhang et al. (2024); Lu et al. (2024), demonstrating the potential for complex world understanding and representing a significant milestone toward the realization of artificial general intelligence.
@@
Recent work Chen et al. (2024) construct a benchmark to evaluate the capability of MLLMs in continual instruction tuning (CIT) and reveal that there is a serious catastrophic forgetting phenomenon in MLLMs. In terms of methodology, they propose MoELoRA to mitigate the model’s forgetting of old instructions. Based on this, ModalPrompt Zeng et al. (2024) dynamically selects optimal prompts during inference by jointly leveraging textual and visual features, thereby enhancing model performance. However, we observe that the downstream datasets used by these methods partially overlaps with the tasks encountered during the supervised fine-tuning (SFT) phase of MLLM. Such information leakage Kim et al. (2023) compromises the reliability of evaluating whether a method mitigates forgetting or if the model inherently retains this capability, thereby diminishing the challenge of continual instruction tuning.
@@ 2 Relate Work
Multimodal Large Language Models. With the predominant capability of large language models (LLMs) Touvron et al. (2023); Zhang et al. (2023a) in handing natural language processing tasks Wang et al. (2022a); Min et al. (2023), huge amount of efforts has been made to multimodal learning and multimodal large language models (MLLMs) Bai et al. (2023); Zhu et al. (2024); Dai et al. (2023) are proposed to tackle task in multimodal scenarios. In addition to a large language model, which is composed of stacks of transformer blocks Vaswani et al. (2017), MLLMs also incorporates a vision encoder Dosovitskiy et al. (2021) and a feature fusion module to align the cross-modality representations (usually a projection layer Liu et al. (2024b) or cross attention Dai et al. (2023)). Answers are then generated in response to multimodal inputs. The paradigm is of great significance and achieves impressive performance on various downstream tasks Liu et al. (2025); Zhang et al. (2024).
Continual Instruction Tuning. Instruction tuning Ouyang et al. (2022); Zhang et al. (2023b) aims to enpower MLLMs to better understand and follow human instructions. To adapt to dynamically changing instructions in real-world scenarios, MLLMs need the capability to learn current and retain previous learned instructions, which is named continual instruction tuning. In recent studies, Eproj He et al. (2023) first constructs a benchmark in continual instruction tuning settings and reveal that catastrophic forgetting is still observed in MLLMs. However, the benchmark they proposed is insufficient in terms of the number and diversity of tasks and is not evaluated on mainstream MLLM architectures. By contrast, CoIN Chen et al. (2024) establishes a diversity benchmark consisting of 8 crafted datasets and proposed MoELoRA to mitigate catastrophic forgetting on various MLLMs. More recently, ModalPrompt Zeng et al. (2024) further enhances performance in this setting by leveraging image and text modalities to guide the selection of appropriate prompts from the prompt pool. Continual-LLaVA Cao et al. (2024) similarly proposes a pool of LoRAs Hu et al. (2021) to select the appropriate LoRAs for training based on textual similarity, and using them during inference.
@@ 4.1 Hierarchical decoupling of MLLM
Existing research Kornblith et al. (2019) shows that the outputs between different layers of the model exhibit notable differences, indicating that the model focuses on different patterns of the input at each layer. Inspired by this, we fine-tune LLaVA-v1.5 Liu et al. (2024b) using LoRA on 6 instruction tuning datasets [^1] Visualizations for each task are provided in Appendix D. and analyze the output CKA similarity Kornblith et al. (2019) differences across the same layers between models. As illustrated in left of Fig 1, the similarity of the model outputs exhibits a clear difference between the top and remaining layers. Specifically, for two datasets that differ significantly in both input image styles and textual instructions (IconQA Lu et al. (2021) and Flickr30k Plummer et al. (2017)), we observe that their output similarity differs notably only in the top layer (i.e. the layer closest to the output), while remaining relatively high across the other layers. This suggests that for different tasks, the model primarily learns similar, generalized knowledge in the layers below the top layer, while the top layer focuses more on dissimilar, task-specific information. Therefore, we hypothesize that for continual instruction tuning, the shared information can be obtained by integrating the layers beneath the top layer, while task-specific knowledge in the top layer should be selectively activated.
@@
To further validate our hypothesis, we performe LoRA fine-tuning on two instruction tuning datasets, ImageNet-R Hendrycks et al. (2021) and CLEVR-Math Lindström and Abraham (2022), to obtain their respective LoRA weights. We then apply the following operations to the LoRAs at different layers: (i) applying corresponding LoRAs to all layers, (ii) applying corresponding LoRAs only to the top layer (32nd layer) while merging the rest (1-31 layers), (iii) merging LoRAs across all layers, (iv) applying only the top layer with non-corresponding LoRAs while merging the rest of the layers, and (v) fine-tuning the same set of LoRAs in sequence. As shown in Fig 2, directly merging all LoRA layers for two tasks provides greater mitigation of catastrophic forgetting compared to sequential fine-tuning (i.e. iii and v), but also leads to performance degradation for the respective tasks. However, merging only the layers below the top layer and correctly selecting task-specific LoRAs for the top layer results in further performance improvements. By contrast, selecting non-corresponding LoRAs for the top layer can degrade the model’s performance on the corresponding task (i.e. ii and iv).
@@ 5.1 Experimental Setup
Datasets We train and evaluate the effectiveness of our method on (1) CoIN Chen et al. (2024) benhcmark and (2) our reconstructed benchmark. In particular, CoIN consists of datasets such as VQAv2 Goyal et al. (2017), VizWiz Gurari et al. (2018), ScienceQA Lu et al. (2022), TextVQA Singh et al. (2019), GQA Hudson and Manning (2019), OCR-VQA Mishra et al. (2019), ImageNet Deng et al. (2009) and REC-COCO Kazemzadeh et al. (2014); Mao et al. (2016). A major limitation of CoIN is the overlap between the selected downstream datasets and the dataset used during LLaVA’s pre-training phase, resulting in information leakage Kim et al. (2023) that significantly undermines fair comparisons between different methods. To address this, we screen six datasets that are highly uncorrelated with the pre-training data based on LLaVA’s zero-shot performance (See Zero-shot in Table 1), specifically: ArxivQA Li et al. (2024), CLEVR-Math Lindström and Abraham (2022), IconQA Lu et al. (2021), ImageNet-R Hendrycks et al. (2021), VizWiz-caption Gurari et al. (2018) and Flickr30k Plummer et al. (2015). We term this the Unseen Continual Instruction Tuning (UCIT) benchmark, with additional details provided in Appendix C.
@@
Baseline. We compare our HiDe-LLaVA with traditional continual learning methods and recent continual instruction tuning method. For the former, we implement LwF Li and Hoiem (2017), EWC Kirkpatrick et al. (2017), L2P Wang et al. (2022b) and O-LoRA Wang et al. (2023) within a MLLM architecture, meticulously tuning parameters to ensure reliable and effective results. For the latter, We compare with MoELoRA Chen et al. (2024) to highlight the superiority of our method. The performance of zero-shot and multi-task fine-tuning is also reported to represent the lower and upper bounds for each benchmark.
@@ 5.2 Main Results
Notably, by comparing the Zero-shot and Multi-task performance on the UCIT and CoIN benchmarks in Table 1 and 2, we observe serious information leakage in CoIN. Specifically, the average performance difference between Multi-task and Zero-shot in CoIN is only 13.19%, while in our UCIT benchmark, the difference is significantly higher at 43.10%. Additionally, some datasets in CoIN even performed worse than Zero-shot after fine-tuning (e.g., GQA, OCRVQA and VQAv2), suggesting that LLaVA may have encountered similar tasks during pre-training. The continued reuse of these datasets could lead to overfitting, thereby impacting the fairness of comparisons between different methods.
@@ Appendix A Details of the comparison method
Appendix B Ablation study of hyper-parameters
Figure 7: Ablation study of temperature coefficient.
Appendix C Details of UCIT benchmark
@@
Appendix D Visualization of UCIT benchmark
@@
Appendix E Details of evaluation

### v2

@@ Abstract
Instruction tuning is widely used to improve a pre-trained Multimodal Large Language Model (MLLM) by training it on curated task-specific datasets, enabling better comprehension of human instructions. However, it is infeasible to collect all possible instruction datasets simultaneously in real-world scenarios. Thus, enabling MLLM with continual instruction tuning is essential for maintaining their adaptability. However, existing methods often trade off memory efficiency for performance gains, significantly compromising overall efficiency. In this paper, we propose a task-specific expansion and task-general fusion framework based on the variations in Centered Kernel Alignment (CKA) similarity across different model layers when trained on diverse datasets. Furthermore, we analyze the information leakage present in the existing benchmark and propose a new and more challenging benchmark to rationally evaluate the performance of different methods. Comprehensive experiments showcase a significant performance improvement of our method compared to existing state-of-the-art methods. Code and dataset are released at https://github.com/Ghy0501/HiDe-LLaVA.
@@
Haiyang Guo1,2thanks: Equal Contribution., Fanhu Zeng2,311footnotemark: 1, Ziwei Xiang2,3, Fei Zhu4, Da-Han Wang5, Xu-Yao Zhang1,2,3thanks: Corresponding Author., Cheng-Lin Liu1,2,3 1School of Advanced Interdisciplinary Sciences, UCAS 2MAIS, CASIA 3School of Artificial Intelligence, UCAS 4Centre for Artificial Intelligence and Robotics, HKISI-CAS 5FKLPRIU, School of Computer and Information Engineering, Xiamen University of Technology {guohaiyang2023, zengfanhu2022, zhufei2018}@ia.ac.cn, {xyz, liucl}@nlpr.ia.ac.cn
@@ 1 Introduction
Recent years have witnessed remarkable advancements in Multimodal Large Language Models (MLLMs) Yin et al. (2023), which extend the capabilities of Large Language Models Touvron et al. (2023) through sophisticated vision-text feature alignment mechanisms Liu et al. (2024b) and autoregressive generation frameworks. The integration of large-scale training corpora and extensive model parameters Yang et al. (2024); Kaplan et al. (2020) has enabled MLLMs to achieve state-of-the-art performance across diverse downstream applications Zhao et al. (2025a); Lu et al. (2024), demonstrating the potential for complex world understanding and representing a significant milestone toward the realization of artificial general intelligence.
@@
Recent work Chen et al. (2024a) construct a benchmark to evaluate the capability of MLLMs in continual instruction tuning (CIT) and reveal that there is a serious catastrophic forgetting phenomenon in MLLMs. In terms of methodology, they propose MoELoRA to mitigate the model’s forgetting of old instructions. Based on this, ModalPrompt Zeng et al. (2024) dynamically selects optimal prompts during inference by jointly leveraging textual and visual features, thereby enhancing model performance. However, we observe that the downstream datasets used by these methods partially overlaps with the tasks encountered during the supervised fine-tuning (SFT) phase of MLLM. Such information leakage Kim et al. (2023) compromises the reliability of evaluating whether a method mitigates forgetting or if the model inherently retains this capability, thereby diminishing the challenge of continual instruction tuning.
@@ 2 Relate Work
Multimodal Large Language Models. With the predominant capability of large language models (LLMs) Touvron et al. (2023); Zhang et al. (2023a) in handing natural language processing tasks Wang et al. (2022a); Min et al. (2023), huge amount of efforts has been made to multimodal learning and multimodal large language models (MLLMs) Bai et al. (2023); Zhu et al. (2024a); Dai et al. (2023) are proposed to tackle task in multimodal scenarios. In addition to a large language model, which is composed of stacks of transformer blocks Vaswani et al. (2017), MLLMs also incorporates a vision encoder Dosovitskiy et al. (2021) and a feature fusion module to align the cross-modality representations (usually a projection layer Liu et al. (2024b) or cross attention Dai et al. (2023)). Answers are then generated in response to multimodal inputs. The paradigm is of great significance and achieves impressive performance on various downstream tasks Liu et al. (2025); Zhao et al. (2025b); Zhang et al. (2024).
Continual Instruction Tuning. Instruction tuning Ouyang et al. (2022); Zhang et al. (2023b) aims to enpower MLLMs to better understand and follow human instructions. To adapt to dynamically changing instructions in real-world scenarios, MLLMs need the capability to learn current and retain previous learned instructions, which is named continual instruction tuning. In recent studies, Eproj He et al. (2023) first constructs a benchmark in continual instruction tuning settings and reveal that catastrophic forgetting is still observed in MLLMs. However, the benchmark they proposed is insufficient in terms of the number and diversity of tasks and is not evaluated on mainstream MLLM architectures. By contrast, CoIN Chen et al. (2024a) establishes a diversity benchmark consisting of 8 crafted datasets and proposed MoELoRA to mitigate catastrophic forgetting on various MLLMs. More recently, ModalPrompt Zeng et al. (2024) further enhances performance in this setting by leveraging image and text modalities to guide the selection of appropriate prompts from the prompt pool. Continual-LLaVA Cao et al. (2024) similarly proposes a pool of LoRAs Hu et al. (2021) to select the appropriate LoRAs for training based on textual similarity, and using them during inference.
@@ 4.1 Hierarchical decoupling of MLLM
Existing research Kornblith et al. (2019) shows that the outputs between different layers of the model exhibit notable differences, indicating that the model focuses on different patterns of the input at each layer. Inspired by this, we fine-tune LLaVA-v1.5 Liu et al. (2024b) using LoRA on 6 instruction tuning datasets and analyze the output CKA similarity Kornblith et al. (2019) differences across the same layers between models. [^1] The details of the CKA similarity computation are presented in Appendix B. As illustrated in left of Fig 1, the similarity of the model outputs exhibits a clear difference between the top and remaining layers. Specifically, for two datasets that differ significantly in both input image styles and textual instructions (IconQA Lu et al. (2021) and Flickr30k Plummer et al. (2017)), we observe that their output similarity differs notably only in the top layer (i.e. the layer closest to the output), while remaining relatively high across the other layers. This suggests that for different tasks, the model primarily learns similar, generalized knowledge in the layers below the top layer, while the top layer focuses more on dissimilar, task-specific information. Therefore, we hypothesize that for continual instruction tuning, the shared information can be obtained by integrating the layers beneath the top layer, while task-specific knowledge in the top layer should be selectively activated.
@@
To further validate our hypothesis, we performe LoRA fine-tuning on two instruction tuning datasets, ImageNet-R Hendrycks et al. (2021) and CLEVR-Math Lindström and Abraham (2022), to obtain their respective LoRA weights. We then apply the following operations to the LoRAs at different layers: (i) applying corresponding LoRAs to all layers, (ii) applying corresponding LoRAs only to the top layer (32-th layer) while merging the rest (1-31 layers), (iii) merging LoRAs across all layers, (iv) applying only the top layer with non-corresponding LoRAs while merging the rest of the layers, and (v) fine-tuning the same set of LoRAs in sequence. As shown in Fig 2, directly merging all LoRA layers for two tasks provides greater mitigation of catastrophic forgetting compared to sequential fine-tuning (i.e. iii and v), but also leads to performance degradation for the respective tasks. However, merging only the layers below the top layer and correctly selecting task-specific LoRAs for the top layer results in further performance improvements. By contrast, selecting non-corresponding LoRAs for the top layer can degrade the model’s performance on the corresponding task (i.e. ii and iv).
@@ 5.1 Experimental Setup
Datasets We train and evaluate the effectiveness of our method on (1) CoIN Chen et al. (2024a) benhcmark and (2) our reconstructed benchmark. In particular, CoIN consists of datasets such as VQAv2 Goyal et al. (2017), VizWiz Gurari et al. (2018), ScienceQA Lu et al. (2022), TextVQA Singh et al. (2019), GQA Hudson and Manning (2019), OCR-VQA Mishra et al. (2019), ImageNet Deng et al. (2009) and REC-COCO Kazemzadeh et al. (2014); Mao et al. (2016). A major limitation of CoIN is the overlap between the selected downstream datasets and the dataset used during LLaVA’s pre-training phase, resulting in information leakage Kim et al. (2023) that significantly undermines fair comparisons between different methods. To address this, we screen six datasets that are highly uncorrelated with the pre-training data based on LLaVA’s zero-shot performance (See Zero-shot in Table 1), specifically: ArxivQA Li et al. (2024), CLEVR-Math Lindström and Abraham (2022), IconQA Lu et al. (2021), ImageNet-R Hendrycks et al. (2021), VizWiz-caption Gurari et al. (2018) and Flickr30k Plummer et al. (2015). We term this the Unseen Continual Instruction Tuning (UCIT) benchmark. Details are provided in Appendix E and F.
@@
Baseline. We compare our HiDe-LLaVA with traditional continual learning methods and recent continual instruction tuning method. For the former, we implement LwF Li and Hoiem (2017), EWC Kirkpatrick et al. (2017), L2P Wang et al. (2022b) and O-LoRA Wang et al. (2023) within a MLLM architecture, meticulously tuning parameters to ensure reliable and effective results. For the latter, We compare with MoELoRA Chen et al. (2024a) to highlight the superiority of our method. The performance of zero-shot and multi-task fine-tuning is also reported to represent the lower and upper bounds for each benchmark. Specific details of each method can be found in Appendix A.
@@ 5.2 Main Results
Notably, by comparing the Zero-shot and Multi-task performance on the UCIT and CoIN benchmarks in Table 1 and 2, we observe serious information leakage in CoIN. Specifically, the average performance difference between Multi-task and Zero-shot in CoIN is only 13.19%, while in our UCIT benchmark, the difference is significantly higher at 43.10%. Additionally, some datasets in CoIN even performed worse than Zero-shot after fine-tuning (e.g., GQA, OCRVQA and VQAv2), suggesting that LLaVA may have encountered similar tasks during pre-training. The continued reuse of these datasets could lead to overfitting, thereby impacting the fairness of comparisons between different methods. Additional experimental results are provided in Appendix D.
@@ 6 Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China (62222609, 62320106010), National Science and Technology Major Project (2022ZD0116500), CAS Project for Young Scientists in Basic Research (YSBR-083), and Key Research Program of Frontier Sciences of CAS (ZDBS-LY-7004), Unveiling and Leading Projects of Xiamen (3502Z20241011) , Major Science and Technology Plan Project on the Future Industry Fields of Xiamen City (3502Z20241027) and the InnoHK program.
@@ Appendix A Details of the comparison method
Figure 7: Ablation study of temperature coefficient.
Appendix B Details of CKA similarity.
The CKA similarity computation process in Section 4.1 can be summarized as:
(1) Assuming that $X_{i}\in\mathbb{R}^{n\times p}$ and $Y_{i}\in\mathbb{R}^{n\times p}$ represent the output features of two different task inputs at layer $i$, where $n$ is the number of samples and $p$ denotes the dimension of the feature. The first step is to compute the linear kernel matrix: $K_{X_{i}}=\hat{X_{i}}\hat{X_{i}}^{T},~{}K_{Y_{i}}=\hat{Y_{i}}\hat{Y_{i}}^{T},$ where $\hat{X_{i}},\hat{Y_{i}}$ denotes the feature after decentralization.
(2) Next, compute the Hilbert-Schmidt Independence Criterion (HSIC) using linear kernel matrices: $\text{HSIC}(K_{X_{i}},K_{Y_{i}})=\frac{1}{(n-1)^{2}}\text{tr}(K_{X_{i}}K_{Y_{i}})$, where $\text{tr}(\cdot)$ denotes the trace of a matrix.
(3) Finally, calculate the CKA similarity by normalizing the HSIC values: $\text{CKA}(X_{i},Y_{i})=\frac{\text{HSIC}(K_{X_{i}},K_{Y_{i}})}{\sqrt{\text{HSIC}(K_{X_{i}},K_{X_{i}})~{}\text{HSIC}(K_{Y_{i}},K_{Y_{i}})}}.$
The value of CKA similarity ranges from 0 to 1, where a higher value indicates greater similarity between the two feature representations. In our work, we compute and compare the CKA similarity for each layer individually using inputs from different datasets.
Appendix C Ablation study of hyper-parameters
Dataset Train Test Task-specific Instruction
ArxivQA 40000 3000 Answer with the option’s letter from the given choices directly.
ImageNet-R 23998 3000 Answer the question using a single word or phrase.
VizWiz-cap 40000 3000 Generate a brief caption for the image.
IconQA 29859 3000 Answer with the option’s letter from the given choices directly.
CLEVR-Math 40000 3000 Answer the question using a single word or phrase.
Flickr30k 40000 3000 Generate a brief caption for the image.
Table 5: Details of datasets used in UCIT benchmark.
@@ Appendix C Ablation study of hyper-parameters
Appendix D Further Experiments
Comparison with Model Tailor. Model Tailor Zhu et al. (2024b) addresses catastrophic forgetting of generic capabilities by identifying and enhancing critical model parameters during downstream task adaptation. However, unlike our setting, which considers continual learning across multiple tasks, Model Tailor only focuses on mitigating forgetting after single-task fine-tuning. To ensure a fair comparison, we preserve Model Tailor’s original methodology while evaluating it on our UCIT benchmark under multi-task continual learning conditions. The comparative results are shown in Table 7. While Model Tailor excels on the first task, its performance declines markedly on subsequent tasks, likely due to its need for repeated parameter recalibration. Our method overcomes this limitation through robust knowledge fusion and expansion, maintaining stable performance across tasks.
      Method R A V I C F Avg
Avg      Model Tailor 89.90 71.30 15.66 39.34 29.50 44.63 48.39
     HiDe-LLaVA 85.70 92.70 54.10 66.87 59.12 55.15 68.94
Last      Model Tailor 88.37 69.00 22.79 38.27 29.60 44.63 48.78
     HiDe-LLaVA 80.50 89.83 48.78 62.90 47.97 55.15 64.19
Table 7: Comparison with Model Tailor on the UCIT benchmark. R: ImageNet-R, A: ArxivQA, V: VizWiz, I: Iconqa, C: CLEVR-Math, F: Flickr30k.
Experiment on other MLLM backbone. To further evaluate the scalability of our proposed method, we conduct experiments on different MLLM architectures. Specifically, we adopt InternVL-Chat-7B Chen et al. (2024b) as our base model, which integrates a 6B-parameter Vision Transformer (ViT) as the visual encoder with a 7B-parameter LLM (i.e. Vicuna) via a multimodal projection layer. The experimental results, presented in Fig. 8, demonstrate consistent performance improvements, confirming our method’s architecture-agnostic scalability.
Figure 8: Experiment on InternVL-Chat-7B.
Appendix E Details of UCIT benchmark
@@
Appendix F Visualization of UCIT benchmark
@@
Appendix G Details of evaluation